# Zero- to low-field -spectroscopy with a diamond magnetometer

### v1

@@ I Introduction
Nuclear Magnetic Resonance (NMR) is a powerful spectroscopic technique that provides detailed information on molecular structure and dynamics, with applications spanning chemistry, biochemistry, materials science, and medicine nmrreview.Traditional high-field NMR relies on strong external magnetic fields to interact with nuclear spins, enabling the detection of signals based on the unique resonances of nuclei in a given chemical environment. However, weak signals, especially in portable benchtop NMR spectrometers with limited field strengths, continue to be a challenge. To address this, various techniques have emerged; for example, parahydrogen induced nuclear polarization (PHIP), which significantly enhances NMR signals through hyperpolarization. We specifically hyperpolarize acetonitrile using signal amplification by reversible exchange (SABRE) BARSKIY2025101558.
@@
Zero-to-ultra-low-field (ZULF) NMR offers a spectroscopic approach by leveraging internal spin interactions rather than relying on strong external fields. The resulting spectral linewidth is narrower than in high-field NMR because of the higher magnetic homogeneity of the shielded environment. This enables high-resolution spectroscopy even of inhomogeneous samples in complex environments BARSKIY2025101558; Kircher2025; Xu2025.
@@
A sensor to be used for the magnetic detection of ZULF NMR spectra of miniature samples ideally needs to be small (¡1 mm) to enable a short stand-off distance, have high sensitivity (in the pT/ range or better), and a wide bandwidth (on the order of 1 kHz). Currently, sensors for ZULF NMR include superconducting quantum interference devices (SQUIDs) Buckenmaier2017, atomic magnetometers LEDBETTER200925, and magnetoresistive (MR) sensors roman. An alternative strategy is shuttling the sample, after evolution under ZULF conditions, into a high magnetic field, allowing for detection by inductive coils BARSKIY2025101558. However, the need for cryogenic cooling limits the usability of SQUIDs. Magnetoresistive sensors, while potentially enabling higher spatial resolution, present challenges due to their residual magnetic fields, which can interfere with the ZULF conditions required for spectral acquisition roman. Optically pumped atomic magnetometers (OPM) can, in principle, achieve high spatial resolution on the order of hundreds of micrometers highspatialresolutionopm, along with sensing bandwidths extending from DC to several hundred kilohertz highbandwidthopm. However, the sensitivity is rapidly lost for small sensor volumes, and the stand-off distance is limited by the vapor cell itself.
@@
This opens an opportunity for nitrogen-vacancy (NV) centers in diamond, which offer superior spatial resolution down to the nanometer scale magneticimagingnanoscale and comparatively straightforward implementation of wide-bandwidth sensing, spanning from DC to beyond megahertz frequencies 2019BarryNVReview. In addition, NV sensors offer several other advantages, including rapid initialization of the electron spin 2019BarryNVReview, robust performance, and the ability to operate at a broad temperature range from cryogenic (350 mK) cryo to far above room temperature (550 K) hotsensing. These sensors have already demonstrated their utility in various areas, for instance, biomagnetic signal detection Barry2016, as well as detection of electric fields electric1, rotations jarmola; PhysRevLett.126.197702 and temperature doherty2014temperature. An overview of the various sensors and their typical characteristics used in ZULF NMR is presented in Tab.1.
@@
| SQUID | 2.5 Matlashov2011 | DC-MHz Myers2024ZeroToUltralow | (coil diameter) | Matlashov et al. Matlashov2011 | |
| Coil | 20 Matlashov2011 | 3 kHz–10 kHz Matlashov2011 | Matlashov2011 (coil diameter) | unspecified | Matlashov et al. Matlashov2011 |
| OPM | LEDBETTER200925 | LEDBETTER200925 | LEDBETTER200925 (cell dimension) | unspecified | Ledbetter et al. LEDBETTER200925 |
| MR | roman | roman | roman | Picazo-Frutos et al. roman |
@@ II NV center zero bias field magnetometry
One of the most common protocols in NV magnetometry is optically detected magnetic resonance (ODMR) 2019BarryNVReview. In this method, the diamond electron spins are polarized by optical pumping into the state in the ground-state electron spin-1 manifold. A microwave (MW) magnetic field is applied to drive transitions within the NV ground-state manifold. These transitions move the NV spins into one or both of the depopulated sublevels, which experience Zeeman frequency shifts under a magnetic field. Transitions can be optically detected 2019BarryNVReview as the photoluminescence is lower for these states compared to . For a single NV center photoluminescence reduces by 30 % 2019BarryNVReview.
@@
Several techniques have been developed to enable zero-field sensing with NV centers. One approach uses circularly polarized MW fields Zheng2020; Lenz2021ZeroFieldNV to selectively excite one of the degenerate sublevels. However, this method is technically challenging because of the difficulty of generating circular polarization along all four NV axes and achieving uniform circularity across the entire diamond volume. Another technique leverages cross-relaxation features near zero magnetic field, which are prominent in NV centers without preferential orientation omkarzero. However, this method has not yet reached sensitivities comparable to standard finite-field ODMR. Magnetically sensitive noise floors have been demonstrated on the order of around 100 nT/ with 2 nT/ photon shot noise Dhungel2024_Near_zero. The method still requires the application of a DC or modulated bias field. Hyperfine coupling with 13C nuclei has also been used for zero-field sensing in single NV centers alternativezerofield. However, extending this to NV ensembles requires the use of isotopically enriched 13C diamond, which increases electron spin dephasing and degrades sensitivity. Furthermore, ensemble averaging over varying NV–13C distances introduces inhomogeneous effective fields that complicate interpretation.
@@
In this study, we employ a zero-field magnetometry protocol with a modulated bias (6.12 kHz along [100] crystallographic direction with T amplitude) achieving sensitivities of 13 pT/ comparable to magnetically-biased ODMR substractionodmr; Zheng2020. The presence of a T-scale magnetic field broadens the zero-field ODMR feature, which, in turn, reduces contrast. This change in peak amplitude can itself be viewed as a form of magnetic resonance, with a peak at zero bias field utilizing all NV axes. Therefore, by modulating and demodulating the magnetic field, a magnetic dispersive signal can be extracted. This signal features a zero crossing at zero bias, allowing for high-sensitivity sensing with a linear response in this regime. This method was already demonstrated and described analytically as an alternative to zero-field sensing using circularly polarized microwaves Zheng2020.
@@
While the zero-field feature frequency is independent to magnetic field changes around zero bias-magnetic fields, temperature variations still result in frequency shifts temperature2. To avoid systematic effects due to these shifts, we stabilize the MW to the zero-field feature. For this we generate an error signal by sinusoidally frequency modulating the MW frequency (modulation frequency 140 kHz, modulation amplitude 100 kHz) and demodulating it with a lock-in amplifier. The demodulated signal is linear to frequency shifts due to temperature variation and just quadratic to frequency shifts due to magnetic field changes. Stabilizing the MW frequency via a proportional-integral-differential (PID) control loop makes the system largely insensitive to temperature fluctuations, enabling long-term measurement campaigns.
@@ III.1 Hyperpolarization
Samples were prepared of ([Ir(IMes)(COD)Cl] (IMes = 1,3-bis(2,4,6-trimethylphenyl)imidazol-2-ylidene, COD = cyclooctadiene), 5 mM) dissolved in acetonitrile with a stabilizing coligand benzylamine (125 mM). The precursor complex was transformed into the SABRE-active polarization transfer catalyst by dissolving parahydrogen (pH2) at a pressure of 7 bar 15 minutes prior to detection Kircher2025. The liquid sample was placed in a spherical glass tube, see Fig. 2, and pH2 was bubbled via a glass capillary with a mass flow of 20 scc min-1. The pH2 gas handling setup used in this work was described previously.Sheberstov2025Roboarms; Kircher2025; Xu2025 Acetonitrile solvent was enriched with [15N]-acetonitrile to 25 % (natural isotopic abundance of 15N nuclei in acetonitrile is around 0.4 % and was increased with pure [15N]-acetonitrile). Hyperpolarization of [15N]-acetonitrile can be generated effectively in ZULF conditions via various protocols, in our case a so-called intermittent bubbling sequence Xu2025. The sequence is sketched in Fig. 1 c).
ZULF NMR signals originate from [15N]-acetonitrile molecules. A pulse of 860 s with 6 T amplitude is applied to initiate the free magnetization decay. The diamond sensor data acquisition lasted 40 s, initiated by a software trigger synchronized with the ZULF sequence. All chemicals except for the catalyst were acquired through Sigma-Aldrich. The catalyst was prepared based on previously developed methods Vazquez2006; Savka2015; Blanchard2021.
@@ III.2 Diamond sensor
The principle of the sensor is based on the ODMR technique 2019BarryNVReview. The sensor head is designed to optimize light collection diamondoptics, see Fig. 2. It includes an output coupler for the optical fiber delivering the pump laser light, and producing a beam of 300 m in diameter. This beam passes through a polarizing beam splitter to clean its polarization. A beam sampler captures part of the light before the beam propagates through the backside of a dichroic mirror and into a lens that focuses the light into the sensing diamond (containing 3.7 ppm of NV and 10 ppm of nitrogen). The sensing diamond is shaped like a truncated pyramid for side collection and mounted on a diamond anvil to enhance the collection efficiency by a factor of four diamondoptics. In addition, a silver reflective coating was applied on the diamond components to avoid leakage of the green laser light and to increase light collection efficiency further. The red photoluminescence (PL) emitted by the diamond is collected with the same lens that focuses the laser beam. This light is reflected by the front side of the dichroic mirror and directed to a photodiode (PD). The sensor uses non-magnetic Hamamatsu photodiodes (S13228-01) soldered onto a printed circuit board (PCB). The current from the green pickup beam PD and the red light PD are subtracted directly on the PCB for differential detection to reduce the sensitivity to laser intensity variation. A 3D-printed clipping mask is used to partially cover the green pick up beam PD , allowing fine adjustment of the differential detection scheme substractionodmr. With 150 mW of green laser light power focused on the diamond, 5 mW of red photoluminescent light is collected on the photodiode. The mobile sensor setup includes an instrument rack on wheels. This rack includes a multi-frequency lock-in amplifier (LIA), a breadboard with a Verdi G5 laser (532 nm) coupled into a high-power single-mode optical fiber that delivers light to the sensor head, a preamplifier to enhance the signal from the sensor head, and a microwave signal generator. The output of the microwave generator is mixed with that of an RF source and then amplified by a microwave amplifier. This frequency-mixing technique alters the hyperfine-resolved spectrum by simultaneously driving two resonance transitions with two microwave frequencies and therefore increases the optical contrast of the electron spin resonance. This in turn directly enhances the magnetic measurement sensitivity 2019BarryNVReview. The microwaves are delivered to the sensor via a flexible PCB. The differential signal of the sensor head is sent to the preamplifier and subsequently to the lock-in amplifier. Around the diamond, a coil is mounted, and its current is modulated with one of the lock-in amplifier outputs to enable the magnetic sensing protocol. The custom designed PCBs for the photodiode and microwave delivery are discussed in the Appendix.
@@
In order to quantify the sensitivity of the diamond sensor we acquire 9 s data traces. The corresponding amplitude spectrum needs to be rescaled to an amplitude spectral density to characterize the noise floor. This requires division by the square root of the effective noise bandwidth (ENBW) fft, in this case, 0.17 Hz considering a Hann window and 9 s duration. We record magnetically sensitive, insensitive and electronic noise spectra. The recorded magnetically sensitive spectral densities are acquired while magnetic field modulation and light are on, the insensitive spectral densities while the magnetic field modulation is off and light is on and the electronic noise spectral densities while both light and magnetic modulation are turned off. The resulting characterization noise spectral densities are shown in Fig. 4 i). The magnetically sensitive noise floor features a 1/f-noise profile and matches the magnetically insensitive noise floor above about 5 Hz. The average of the magnetically insensitive noise spectral density below 50 Hz is 13 pT/.
@@ III.3 The vapor cell magnetometer
The optically pumped magnetometer is a commercially available hot vapor cell based sensor (QUSPIN) employing rubidium atoms to measure the magnetic-field-dependent absorption of laser lightquspin; zerofieldresonance. The bandwidth of the sensor is specified as 150 Hz, the dynamic range as nT and the sensitivity on the order of 10 fT/ quspin. The OPM has a sharp hardware digital low-pass filter at 500 Hzquspin which inhibits sensing fields beyond that frequency. The OPM signal was recorded after initialization by its commercial software and read-out with a custom-made python code. The OPM data is acquired without averaging.

### v2

@@ I Introduction
Nuclear Magnetic Resonance (NMR) is a powerful spectroscopic technique that provides detailed information on molecular structure and dynamics, with applications spanning chemistry, biochemistry, materials science, and medicine Wong (2014).Traditional high-field NMR relies on strong external magnetic fields to interact with nuclear spins, enabling the detection of signals based on the unique resonances of nuclei in a given chemical environment. However, weak signals, especially in portable benchtop NMR spectrometers with limited field strengths, continue to be a challenge. To address this, various techniques have emerged; for example, parahydrogen induced nuclear polarization (PHIP), which significantly enhances NMR signals through hyperpolarization. We specifically hyperpolarize acetonitrile using signal amplification by reversible exchange (SABRE) Barskiy et al. (2025).
@@
Zero-to-ultra-low-field (ZULF) NMR offers a spectroscopic approach by leveraging internal spin interactions rather than relying on strong external fields. The resulting spectral linewidth is narrower than in high-field NMR because of the higher magnetic homogeneity of the shielded environment. This enables high-resolution spectroscopy even of inhomogeneous samples in complex environments Barskiy et al. (2025); Kircher et al. (2025); Xu,J et al. (2025).
@@
A sensor to be used for the magnetic detection of ZULF NMR spectra of miniature samples ideally needs to be small (¡1 mm) to enable a short stand-off distance, have high sensitivity (in the pT/ range or better), and a wide bandwidth (on the order of 1 kHz). Currently, sensors for ZULF NMR include superconducting quantum interference devices (SQUIDs) Buckenmaier et al. (2017), atomic magnetometers Ledbetter et al. (2009), and magnetoresistive (MR) sensors Picazo-Frutos,Román et al. (2025). An alternative strategy is shuttling the sample, after evolution under ZULF conditions, into a high magnetic field, allowing for detection by inductive coils Barskiy et al. (2025). However, the need for cryogenic cooling limits the usability of SQUIDs. Magnetoresistive sensors, while potentially enabling higher spatial resolution, present challenges due to their residual magnetic fields, which can interfere with the ZULF conditions required for spectral acquisition Picazo-Frutos,Román et al. (2025). Optically pumped atomic magnetometers (OPM) can, in principle, achieve high spatial resolution on the order of hundreds of micrometers Fang,Xiujie et al. (2020), along with sensing bandwidths extending from DC to several hundred kilohertz Li et al. (2020). However, the sensitivity is rapidly lost for small sensor volumes, and the stand-off distance is limited by the vapor cell itself.
@@
This opens an opportunity for nitrogen-vacancy (NV) centers in diamond, which offer superior spatial resolution down to the nanometer scale Lee et al. and comparatively straightforward implementation of wide-bandwidth sensing, spanning from DC to beyond megahertz frequencies Barry et al. (2020). In addition, NV sensors offer several other advantages, including rapid initialization of the electron spin Barry et al. (2020), robust performance, and the ability to operate at a broad temperature range from cryogenic (350 mK) Scheidegger et al. (2022) to far above room temperature (550 K) Attrash,Mohammed et al. (2023). These sensors have already demonstrated their utility in various areas, for instance, biomagnetic signal detection Barry et al. (2016), as well as detection of electric fields Dolde et al. , rotations Jarmola (2021); Soshenko et al. (2021) and temperature Doherty et al. (2014). An overview of the various sensors and their typical characteristics used in ZULF NMR is presented in Tab.1.
@@
| SQUID | 2.5 Matlashov et al. (2011) | DC-MHz Myers et al. (2024) | (coil diameter) | Matlashov et al. Matlashov et al. (2011) | |
| Coil | 20 Matlashov et al. (2011) | 3 kHz–10 kHz Matlashov et al. (2011) | Matlashov et al. (2011) (coil diameter) | unspecified | Matlashov et al. Matlashov et al. (2011) |
| OPM | Ledbetter et al. (2009) | Ledbetter et al. (2009) | Ledbetter et al. (2009) (cell dimension) | unspecified | Ledbetter et al. Ledbetter et al. (2009) |
| MR | Picazo-Frutos,Román et al. (2025) | Picazo-Frutos,Román et al. (2025) | Picazo-Frutos,Román et al. (2025) | Picazo-Frutos et al. Picazo-Frutos,Román et al. (2025) |
@@ II NV center zero bias field magnetometry
One of the most common protocols in NV magnetometry is optically detected magnetic resonance (ODMR) Barry et al. (2020). In this method, the diamond electron spins are polarized by optical pumping into the state in the ground-state electron spin-1 manifold. A microwave (MW) magnetic field is applied to drive transitions within the NV ground-state manifold. These transitions move the NV spins into one or both of the depopulated sublevels, which experience Zeeman frequency shifts under a magnetic field. Transitions can be optically detected Barry et al. (2020) as the photoluminescence is lower for these states compared to . For a single NV center photoluminescence reduces by 30 % Barry et al. (2020).
@@
Several techniques have been developed to enable zero-field sensing with NV centers. One approach uses circularly polarized MW fields Zheng et al. (2020); Lenz et al. (2021) to selectively excite one of the degenerate sublevels. However, this method is technically challenging because of the difficulty of generating circular polarization along all four NV axes and achieving uniform circularity across the entire diamond volume. Another technique leverages cross-relaxation features near zero magnetic field, which are prominent in NV centers without preferential orientation Dhungel,Omkar et al. (2024). However, this method has not yet reached sensitivities comparable to standard finite-field ODMR. Magnetically sensitive noise floors have been demonstrated on the order of around 100 nT/ with 2 nT/ photon shot noise Dhungel et al. (2024). The method still requires the application of a DC or modulated bias field. Hyperfine coupling with 13C nuclei has also been used for zero-field sensing in single NV centers Wang,Ning et al. (2022). However, extending this to NV ensembles requires the use of isotopically enriched 13C diamond, which increases electron spin dephasing and degrades sensitivity. Furthermore, ensemble averaging over varying NV–13C distances introduces inhomogeneous effective fields that complicate interpretation.
@@
In this study, we employ a zero-field magnetometry protocol with a modulated bias (6.12 kHz along [100] crystallographic direction with T amplitude) achieving sensitivities of 13 pT/ comparable to magnetically-biased ODMR Sekiguchi,N. et al. (2024); Zheng et al. (2020). The presence of a T-scale magnetic field broadens the zero-field ODMR feature, which, in turn, reduces contrast. This change in peak amplitude can itself be viewed as a form of magnetic resonance, with a peak at zero bias field utilizing all NV axes. Therefore, by modulating and demodulating the magnetic field, a magnetic dispersive signal can be extracted. This signal features a zero crossing at zero bias, allowing for high-sensitivity sensing with a linear response in this regime. This method was already demonstrated and described analytically as an alternative to zero-field sensing using circularly polarized microwaves Zheng et al. (2020).
@@
While the zero-field feature frequency is independent to magnetic field changes around zero bias-magnetic fields, temperature variations still result in frequency shifts Shim et al. . To avoid systematic effects due to these shifts, we stabilize the MW to the zero-field feature. For this we generate an error signal by sinusoidally frequency modulating the MW frequency (modulation frequency 140 kHz, modulation amplitude 100 kHz) and demodulating it with a lock-in amplifier. The demodulated signal is linear to frequency shifts due to temperature variation and just quadratic to frequency shifts due to magnetic field changes. Stabilizing the MW frequency via a proportional-integral-differential (PID) control loop makes the system largely insensitive to temperature fluctuations, enabling long-term measurement campaigns.
@@ III.1 Hyperpolarization
Samples were prepared of ([Ir(IMes)(COD)Cl] (IMes = 1,3-bis(2,4,6-trimethylphenyl)imidazol-2-ylidene, COD = cyclooctadiene), 5 mM) dissolved in acetonitrile with a stabilizing coligand benzylamine (125 mM). The precursor complex was transformed into the SABRE-active polarization transfer catalyst by dissolving parahydrogen (pH2) at a pressure of 7 bar 15 minutes prior to detection Kircher et al. (2025). The liquid sample was placed in a spherical glass tube, see Fig. 2, and pH2 was bubbled via a glass capillary with a mass flow of 20 scc min-1. The pH2 gas handling setup used in this work was described previously.Sheberstov et al. (2025); Kircher et al. (2025); Xu,J et al. (2025) Acetonitrile solvent was enriched with [15N]-acetonitrile to 25 % (natural isotopic abundance of 15N nuclei in acetonitrile is around 0.4 % and was increased with pure [15N]-acetonitrile). Hyperpolarization of [15N]-acetonitrile can be generated effectively in ZULF conditions via various protocols, in our case a so-called intermittent bubbling sequence Xu,J et al. (2025). The sequence is sketched in Fig. 1 c).
ZULF NMR signals originate from [15N]-acetonitrile molecules. A pulse of 860 s with 6 T amplitude is applied to initiate the free magnetization decay. The diamond sensor data acquisition lasted 40 s, initiated by a software trigger synchronized with the ZULF sequence. All chemicals except for the catalyst were acquired through Sigma-Aldrich. The catalyst was prepared based on previously developed methods Vazquez-Serrano et al. (2006); Savka and Plenio (2015); Blanchard et al. (2021b).
@@ III.2 Diamond sensor
The principle of the sensor is based on the ODMR technique Barry et al. (2020). The sensor head is designed to optimize light collection Omar,Muhib et al. (2023), see Fig. 2. It includes an output coupler for the optical fiber delivering the pump laser light, and producing a beam of 300 m in diameter. This beam passes through a polarizing beam splitter to clean its polarization. A beam sampler captures part of the light before the beam propagates through the backside of a dichroic mirror and into a lens that focuses the light into the sensing diamond (containing 3.7 ppm of NV and 10 ppm of nitrogen). The sensing diamond is shaped like a truncated pyramid for side collection and mounted on a diamond anvil to enhance the collection efficiency by a factor of four Omar,Muhib et al. (2023). In addition, a silver reflective coating was applied on the diamond components to avoid leakage of the green laser light and to increase light collection efficiency further. The red photoluminescence (PL) emitted by the diamond is collected with the same lens that focuses the laser beam. This light is reflected by the front side of the dichroic mirror and directed to a photodiode (PD). The sensor uses non-magnetic Hamamatsu photodiodes (S13228-01) soldered onto a printed circuit board (PCB). The current from the green pickup beam PD and the red light PD are subtracted directly on the PCB for differential detection to reduce the sensitivity to laser intensity variation. A 3D-printed clipping mask is used to partially cover the green pick up beam PD , allowing fine adjustment of the differential detection scheme Sekiguchi,N. et al. (2024). With 150 mW of green laser light power focused on the diamond, 5 mW of red photoluminescent light is collected on the photodiode. The mobile sensor setup includes an instrument rack on wheels. This rack includes a multi-frequency lock-in amplifier (LIA), a breadboard with a Verdi G5 laser (532 nm) coupled into a high-power single-mode optical fiber that delivers light to the sensor head, a preamplifier to enhance the signal from the sensor head, and a microwave signal generator. The output of the microwave generator is mixed with that of an RF source and then amplified by a microwave amplifier. This frequency-mixing technique alters the hyperfine-resolved spectrum by simultaneously driving two resonance transitions with two microwave frequencies and therefore increases the optical contrast of the electron spin resonance. This in turn directly enhances the magnetic measurement sensitivity Barry et al. (2020). The microwaves are delivered to the sensor via a flexible PCB. The differential signal of the sensor head is sent to the preamplifier and subsequently to the lock-in amplifier. Around the diamond, a coil is mounted, and its current is modulated with one of the lock-in amplifier outputs to enable the magnetic sensing protocol. The custom designed PCBs for the photodiode and microwave delivery are discussed in the Appendix.
@@
In order to quantify the sensitivity of the diamond sensor we acquire 9 s data traces. The corresponding amplitude spectrum needs to be rescaled to an amplitude spectral density to characterize the noise floor. This requires division by the square root of the effective noise bandwidth (ENBW) Heinzel et al. (2002), in this case, 0.17 Hz considering a Hann window and 9 s duration. We record magnetically sensitive, insensitive and electronic noise spectra. The recorded magnetically sensitive spectral densities are acquired while magnetic field modulation and light are on, the insensitive spectral densities while the magnetic field modulation is off and light is on and the electronic noise spectral densities while both light and magnetic modulation are turned off. The resulting characterization noise spectral densities are shown in Fig. 4 i). The magnetically sensitive noise floor features a 1/f-noise profile and matches the magnetically insensitive noise floor above about 5 Hz. The average of the magnetically insensitive noise spectral density below 50 Hz is 13 pT/.
@@ III.3 The vapor cell magnetometer
The optically pumped magnetometer is a commercially available hot vapor cell based sensor (QUSPIN) employing rubidium atoms to measure the magnetic-field-dependent absorption of laser lightM. Zhao and Fujimoto ; Dupont-Roc et al. . The bandwidth of the sensor is specified as 150 Hz, the dynamic range as nT and the sensitivity on the order of 10 fT/ M. Zhao and Fujimoto . The OPM has a sharp hardware digital low-pass filter at 500 HzM. Zhao and Fujimoto which inhibits sensing fields beyond that frequency. The OPM signal was recorded after initialization by its commercial software and read-out with a custom-made python code. The OPM data is acquired without averaging.